### v1

# HiLo: Detailed and Robust 3D Clothed Human Reconstruction with High-and Low-Frequency Information of Parametric Models

Abstract
Reconstructing 3D clothed human involves creating a detailed geometry of individuals in clothing, with applications ranging from virtual try-on, movies, to games. To enable practical and widespread applications, recent advances propose to generate a clothed human from an RGB image. However, they struggle to reconstruct detailed and robust avatars simultaneously. We empirically find that the high-frequency (HF) and low-frequency (LF) information from a parametric model has the potential to enhance geometry details and improve robustness to noise, respectively. Based on this, we propose HiLo, namely clothed human reconstruction with high- and low-frequency information, which contains two components. 1) To recover detailed geometry using HF information, we propose a progressive HF Signed Distance Function to enhance the detailed 3D geometry of a clothed human. We analyze that our progressive learning manner alleviates large gradients that hinder model convergence. 2) To achieve robust reconstruction against inaccurate estimation of the parametric model by using LF information, we propose a spatial interaction implicit function. This function effectively exploits the complementary spatial information from a low-resolution voxel grid of the parametric model. Experimental results demonstrate that HiLo outperforms the state-of-the-art methods by $10.43\%$ and $9.54\%$ in terms of Chamfer distance on the Thuman2.0 and CAPE datasets, respectively. Additionally, HiLo demonstrates robustness to noise from the parametric model, challenging poses, and various clothing styles.  [^1] Code link: https://github.com/YifYang993/HiLo.git
1 Introduction
Figure 1: Visualization comparisons on in-the-wild images, our HiLo achieves more accurate and detailed reconstruction on challenging poses and diverse clothes.
The creation of 3D realistic digital human plays a pivotal role in the realm of mixed reality [23, 53, 33], remote presentation [53, 5], film [46, 16], and gaming [48]. Traditional methods often require expensive and specialized equipment combined with complex artistic efforts to customize the avatars[43, 17, 58], which limits the ability of individuals to create personalized avatars easily. To address the limitation, recent approaches [1, 2, 28, 41, 42, 31, 25, 60, 50, 51, 52] capture a 3D avatar from an RGB image of a clothed human, thus eliminating the need for costly scanning equipment and making it easier for a broader range of users to create personalized avatars.
Despite the convenience of recent advances, the input image usually lacks details about delicate human body parts and diverse clothes from multiple angles. Moreover, the limited viewpoints and lack of accurate depth information make the reconstruction vulnerable to noise, e.g., inaccurate shape and pose of the estimated parametric body model [54, 55, 15]. Therefore, a detailed and robust 3D human reconstruction from an RGB image is challenging.
In spite of the impressive results of the previous methods, they have not fully addressed the problem of detailed and robust reconstruction simultaneously. Specifically, PIFu [41] produces overly smoothed or non-human body shapes on the unseen side of the human from the input image. ECON [51] requires Poisson Surface Reconstruction and replacement of body parts, introducing an extent of computational overhead (c.f. Sec. 4.4). Additionally, there is a risk of body part misalignment when the mid-term data is inaccurate. Considering that clothes need to conform to the surface of naked bodies, the geometry of the parametric model provides effective semantic regularization for reconstructing clothed humans. PaMIR [60], ICON [50], and D-IF [52] use parametric human bodies  [26, 38] to regularize the reconstruction. However, the performance of these methods degrades significantly when facing noise on the parameters from the estimated naked bodies (c.f. Sec. 4.4).
Figure 2: We empirically demonstrate the effectiveness of the high-frequency (HF) regularization from naked bodies in enhancing geometry details in Toy Experiment Refer to caption. We also verify the effectiveness of the low-frequency (LF) regularization in improving robustness to noise in Toy Experiment Refer to caption.
To achieve robust 3D clothed human reconstruction with detailed geometry, we aim to explore how to further use the regularization from the parametric model to facilitate this goal. Our exploration is based on two common observations. First, high-frequency (HF) information enhances details [40, 30]. Considering that Signed Distance Function (SDF) [34] describes the geometry of a parametric model by representing a distance to the object surface boundary, we investigate the effectiveness of SDF in improving the geometry details of clothed humans. Second, low-frequency (LF) information is relatively robust to noise [10, 11, 24, 57]. Since inaccurate parametric model estimation within an error range has an insignificant impact on the corresponding low-resolution voxel grid [45], we seek to use the voxel grid to mitigate the noise of the estimated body. As shown in Fig. 2, qualitative results demonstrate that SDF boosts the reconstruction details while the voxel grid improves robustness to noise. However, how to effectively combine high- and low-frequency information to generate details and mitigate noise simultaneously is still an open question.
In this paper, we propose a high- and low-frequency paradigm HiLo, which stands for clothed human reconstruction with high- and low-frequency information. To achieve HF detail, we further enhance SDF with HF function [40]. Intuitively, by amplifying the variation of adjacent points that share similar SDFs, we allow for better delineation and capturing of fine details in the 3D human. Moreover, to alleviate the convergence difficulty caused by the large gradients amplified by HF function (c.f. Sec. 3), we introduce a progressive HF SDF that learns detailed 3D geometry in a coarse-to-fine manner. To achieve robustness, we seek to capture the LF complementary information of the low-resolution voxel grid from the naked human body. To this end, we design a spatial interaction implicit function, which promotes the interaction of global and local information across different voxels via an attention mechanism.
We qualitatively and quantitatively evaluate our HiLo on in-the-wild images and benchmark datasets. The experimental results verify the superiority of HiLo over previous approaches in three key aspects: 1) 3D geometry details (see Fig. 1). 2) Robust reconstruction. 3) Convergence speed. We summarize our contributions in three folds:
  • To enhance the geometry details and improve robustness against noise during the clothed human reconstruction process, we propose to explore the high-frequency (HF) information and low-frequency (LF) information from a parametric body model simultaneously.
  • To facilitate detailed reconstruction, we introduce a progressive HF function to enhance the signed distance function (SDF) of a parametric model, providing regularization during the reconstruction process. This function learns an HF SDF in a progressive manner to alleviate the convergence difficulty associated with HF information. Experimental results show that HiLo reconstructs a more detailed clothed human.
  • To ensure robust reconstruction, we employ LF information of low-resolution voxel grids from the parametric model to regularize the reconstruction. We propose a spatial interaction implicit function that reasons complementary information between different voxels. Experimental results show that HiLo is robust to various levels of noise.
2 Preliminaries
2.1 Signed Distance Function
Signed distance function (SDF) [34] is a continuous function that takes a given spatial point $p$ with spatial coordinate $x\in\mathbb{R}^{n}$ and outputs the distance $s\in\mathbb{R}$ of the point to the closest point on the surface $\partial\Omega$ of an object $\Omega$:
$\mathcal{F}_{s}(\mathbf{p})=s,~{}~{}s=\left\{\begin{array}[]{rc}d(x,\partial\Omega)&\mathrm{if}~{}x\not\in\Omega,\\ -d(x,\partial\Omega)&\mathrm{if}~{}x\in\Omega,\\ \end{array}\right.$ (1)
where $d(x,\partial\Omega)=\mathop{\inf}\limits_{y\in\partial\Omega}(||x-y||_{2})$. The sign of the distance implies whether the point is inside (negative) or outside (positive) of the surface $\partial\Omega$. $s=0$ denotes the point $\mathbf{p}$ locates on the $\partial\Omega$.
2.2 Voxel Grid and Mesh Voxelization
Voxel Grids $\Omega_{v}\in\mathbb{R}^{d\times h\times w}$ is a representative data structure for describing a 3D object $\Omega$. Specifically, $\Omega_{v}$ is a three-dimensional matrix of 3D space with depth $d$, height $h$ and width $w$. $\Omega_{v}$ is composed of equally distributed and equally sized cube-shaped volumetric elements called voxels. The term voxel is the 3D counterpart to a 2D pixel. The resolution of a voxel grid is determined by the size of the voxels and the dimensions of the grid. Lower resolution implies larger voxels, resulting in a coarser representation of the space. Thus, we use the low-resolution voxel grid to represent the low-frequency information of a 3D object.
Mesh voxelization $\mathcal{V}$ is a computational technique that plays a crucial role in converting irregular continuous 3D geometric models [4, 26, 38] such as triangular mesh and point clouds into regular and discrete voxel grids. In this process, a 3D mesh $\Omega_{m}$, which is a collection of connected triangles, is converted into a grid of voxels $\Omega_{v}\in\mathbb{R}^{3}$. In this paper, we use $\mathcal{V}$ to transfer SMPL-X mesh $\mathcal{M}$ to a low-resolution voxelized mesh $\mathcal{M}_{v}\in\mathbb{R}^{32\times 32\times 32}$, and then use a 3D CNN to encode $\mathcal{M}_{v}$ for $\mathcal{M}_{v}^{3D}\in$ for more flexibility. Based on our observation from Tab. 3, the low-frequency natural of $\mathcal{M}_{v}^{3D}$ aids multiple existing methods in mitigating various noise levels in SMPL-X shape and pose.
3 Clothed Human Reconstruction with High- and Low-frequency Information
Figure 3: Overview of our proposed HiLo. Conditioned on a single-view image $\mathcal{I}$ and the corresponding SMPL-X $\mathcal{M}$, we first prepare a signed distance field $s$ and a low-resolution voxel grid $\mathcal{M}_{v}^{3D}$ of the naked body. Then, our proposed progressive high-frequency signed distance function $\mathcal{H}(s;\beta)$ enhances $s$ for detailed geometry of the clothed human and alleviates convergence difficulties introduced by large gradients in a coarse-to-fine learning manner. Moreover, we design an implicit function $\phi_{si}$ which leverages the complementary information of low-frequency voxels from $\mathcal{M}_{v}^{3D}$ to mitigate various levels of noise. Finally, we combine the above HF and LF features to $\phi_{si}$ to infer the occupancy field $\mathcal{\hat{O}}$ of the clothed avatar.
We aim to robustly infer detailed 3D clothed avatars from RGB images $\mathcal{I}$. Recent advances [8, 9, 20, 19] tend to use parametric naked body $\mathcal{M}$ such as SMPL [26] or SMPL-X [38] estimated from $\mathcal{I}$ to provide semantic regularization on clothed human avatars. We empirically verify that high-frequency (HF) and low-frequency (LF) information from $\mathcal{M}$ are able to refine geometry and improve robustness to the reconstruction of the clothed human (Sec. 4.1). Based on this, we propose clothed human reconstruction with high- and low-frequency information, namely HiLo, which balances the HF and LF information to achieve detailed and robust reconstruction simultaneously. As shown in Fig. 3, HiLo contains two key components: (1) To refine the geometry of clothed human with HF information, we propose to use progressive high-frequency function to enhance the signed distance function (SDF) of $\mathcal{M}$ (c.f. Sec. 2), and alleviate the convergence difficulty introduced by large gradients in a coarse-to-fine enhancement manner. (2) To achieve robust reconstruction using low-frequency information, we explore complementary information from low-resolution voxels $\mathcal{M}_{v}$ from $\mathcal{M}$ for a more comprehensive understanding of human geometry. To this end, we design a spatial interactive implicit function (c.f. Sec. 3.2) that leverages spatial information from local and global voxelized SMPL-X to predict the occupancy field $\mathcal{\hat{O}}$. Finally, we use the Marching Cubes ([27]) to obtain the 3D mesh of the clothed avatar from $\mathcal{\hat{O}}$.
The overall optimization of our proposed HiLo minimizes the following objective function:
$\mathcal{L}_{overall}=\mathcal{L}_{a}(\mathcal{\hat{O}},\mathcal{O}),$ (2)
where $\mathcal{L}_{a}$ is the MSE loss and $\mathcal{O}$ is a GT occupancy field.
3.1 Progressive Growing High-Frequency SDF
Given that SDFs can enhance 3D reconstruction quality as confirmed in Sec. 4.1, we will leverage this for more realistic avatar reconstruction. However, directly fitting input coordinates with SDFs may lead to subpar representation of high-frequency variation in geometry (see Sec. 4.3). This aligns with previous work [39] indicating neural networks prioritize learning low-frequency signals. We will explore effective strategies to mitigate this.
Conventional high-frequency SDF. To improve the ability to represent complicated 3D shapes robustly, a straightforward way is to apply periodic functions $\mathcal{H}$ such as sine and cosine [44] to extract high-frequency signals on SDF of each sampled point via
$\mathcal{H}(s)$ $=[s,\mathcal{H}_{0}(s),\mathcal{H}_{1}(s),\ldots,\mathcal{H}_{k}(s),\ldots,\mathcal{H}_{L}(s)],$ (3)
$\mathcal{H}_{k}(s)$ $=[\sin(2^{k}\pi s),\cos(2^{k}\pi s)],k\in\{0,1,\ldots,L\}.$
In this way, we amplify the variation of adjacent points that share similar SDFs, allowing for better delineation and capturing of fine details in the 3D object.
Despite the positive characteristic of high-frequency SDF, effective updating for parameters is difficult. Specifically, the gradient of $\mathcal{H}_{k}(s)$ w.r.t. $s$ is calculated by
$\frac{\partial\mathcal{H}_{k}(s)}{\partial s}=2^{k}\pi[\cos{(2^{k}\pi s)},-\sin{(2^{k}\pi s)}].$ (4)
Eqn. (4) incorporated with the coefficient $2^{k}\pi$ will significantly amplify the gradient signals regarding SDF, especially for larger $k$. Large gradients could lead to convergence difficulties and numerical instability, ultimately resulting in poor representation performance [36, 3].
High-frequency SDF in a growing manner. To address the above issue, we introduce a progressively growing approach as shown in Fig. 3 (2), initially emphasizing low-frequency signal learning and gradually focusing on learning the high-frequency geometry. Specifically, in the early stage of training, we reduce the weight of high-frequency signals (e.g.,$\mathcal{H}_{k}(s)$) which have higher $k$ and progressively increase their importance during training. We formulate this schedule as $\mathcal{H}_{k}(s;\beta)$ with a weight $\omega_{k}(\beta)$ following [35]:
$\mathcal{H}_{k}(s;\beta)=\omega_{k}(\beta)[\sin{(2^{k}\pi s)},\cos{(2^{k}\pi s)}],$ (5)
$\omega_{k}(\beta)=\left\{\begin{array}[]{ll}0&\mathrm{if}~{}\beta-k<0;\\ \frac{1-\cos((\beta-k)\pi)}{2}&\mathrm{if}~{}0\leq\beta-k<1;\\ 1&\mathrm{if}~{}\beta-k>1,\\ \end{array}\right.$
where $\beta$ is proportional to the iteration of the optimization process, see Fig. 4 for the relationship between $\omega_{k}(\beta)$ and $\beta$. With $\omega_{k}(\beta)$, the gradient of $\mathcal{H}_{k}(s;\beta)$ becomes
$\frac{\partial\mathcal{H}_{k}(s;\beta)}{\partial s}=\omega_{k}(\beta)2^{k}\pi[\cos{(2^{k}\pi s)},-\sin{(2^{k}\pi s)}].$ (6)
Then, during the beginning of the optimization, $\beta$ is set so small that only frequency components with a smaller value of $k$ will be assigned a non-zero weight, while the frequency components with a higher value of $k$ will be omitted. Throughout the optimization, the higher-frequency components are progressively activated. This manner allows HiLo to explore the low-frequency part and later focus on the fine-grained geometry of 3D humans.
Figure 4: Illustration of the relationship between progressive weights $\omega$ and $\beta$ during the training process.
3.2 Low-Frequency Information for Robustness
Most recent methods  [51, 50, 52, 60] are based on SMPL-X. However, SMPL-X estimation often faces misalignment issues with the corresponding image, especially when facing a challenging human pose. Thus, it is crucial to achieve robust reconstruction against misaligned SMPL-X. Our results (c.f. Fig. 2) show that the low-frequency information, represented by low-resolution voxel grids $\mathcal{M}_{v}$ of SMPL-X $\mathcal{M}$, enhances robustness against noise. We leverage this insight to incorporate local and global information of $\mathcal{M}_{v}$ for improved regularization in reconstruction.
Local voxels for 3D feature preparation. Motivated by that point-wise local 3D features from $\mathcal{M}_{v}$ are robust to out-of-distribution pose and shape [60], we voxelize the estimated SMPL-X $\mathcal{M}$ and query it by $\mathbf{p}$. Specifically, to obtain the voxelization features, we convert the corresponding SMPL-X $\mathcal{M}$ into a low-resolution voxel grid $\mathcal{M}_{\mathcal{V}}$ by mesh voxelization operation $\mathcal{V}$ [60] and encode it via a 3D CNN $f_{\text{3D}}$ for a 3D feature volume $\mathcal{M}^{3D}_{\mathcal{V}}$. To obtain point-wise 3D features, we use trilinear interpolation to sample $\mathcal{M}_{v}^{3D}$ based on coordinate $\mathbf{p}$ of sampled 3D points, resulting in $\mathcal{M}^{3D}_{\mathcal{V}}(\mathbf{p})$. We empirically find that by combining $\mathcal{M}^{3D}_{\mathcal{V}}(\mathbf{p})$, HiLo is robust to SMPL-X noise (c.f. Sec. 4.4). As shown in Fig. 3 (1), in addition to $\mathcal{H}(s;\beta)$ and $\mathcal{M}^{3D}_{\mathcal{V}}$, we follow ICON [50] to use a normal feature $\mathbf{F}_{\mathrm{n}}(\mathbf{p})$ to provide detailed texture information. Then, we concatenate them into one final feature $\mathbf{F}_{c}^{1}$=$[\mathcal{H}(s;\beta),\mathcal{M}^{3D}_{\mathcal{V}}({\mathbf{p}}),\mathbf{F}_{\mathrm{n}}(\mathbf{p})]$ and then fed $\mathbf{F}_{c}^{1}$ to our designed implicit function to reconstruct the clothed avatar.
Global voxels for spatial interaction implicit function. To reconstruct clothed avatars, a typical solution is to map 3D features $\mathbf{F}$ to a continuous occupancy field that represents the interior and exterior of a clothed human. To this end, numerous literature [41, 60, 50, 52] uses an implicit function parameterized by a memory-efficient multi-layer perceptron (MLP) $\mathcal{T}$ to map $\mathbf{F}$ into an occupancy field $\mathcal{\hat{O}}$.
However, the potential issue of the existing implicit function lies in its underutilization of the global information inherent in 3D data. Previous research [6, 32, 21] has shown that the representation ability of features can be improved by capturing the global correlation between the features. For 3D clothed human reconstruction, different human body parts contain distinct yet complementary spatial information. For instance, as shown in Fig. 5 (a), the voxels located on the shoulder (the red point) may offer valuable topological cues to constrain the prediction of geometry near the elbow (the blue point).
Figure 5: (a) Complementarity of voxel Refer to caption and Refer to caption. (b) Illustration of the spatial interaction module $\mathcal{A}$.
To leverage global information from the voxel grid of SMPL-X of $\mathcal{M}_{v}^{3D}$, we design a spatial interaction module $\mathcal{A}$ into $\phi$ to infer the 3D occupancy, denoted by $\phi_{si}$, see detail in Appx B.2. As shown in Fig. 5 (b), $\phi_{si}$ injects global-scale features of $\mathcal{M}_{v}^{3D}$ to the local 3D feature with the aim of introducing whole-body awareness to $\phi_{si}$:
${\phi_{si}}(\mathbf{F}_{c}^{1})\to\mathcal{\hat{O}},~{}~{}\phi_{si}(\cdot)=\mathcal{A}^{N+1}\circ T^{(N+1)}\circ\cdots\circ\mathcal{A}^{1}(\cdot)\circ T^{(1)}.$ (7)
Optimization. We optimize parameters of $\phi_{si}$ and $f_{3D}$ via minimizing the MSE loss in Eqn. (2) between the predicted occupancy field $\mathcal{\hat{O}}$ and the ground-truth occupancy field $\mathcal{O}$. With $\mathcal{\hat{O}}$, we reconstruct the triangular mesh of the 3D clothed avatar via marching cubes algorithm  [27].
4 Experiments
Datasets: We conduct experiments on two open-source datasets, i.e., Thuman2.0 [59] and CAPE [28], which both contain various human shapes with different human poses and diverse clothes. Specifically, following ICON [50], the CAPE dataset is divided into the "CAPE-FP" and "CAPE-NFP" sets, which have "fashion" and "non-fashion" poses, respectively, to further analyze the generalization to complex body poses.
Moreover, to evaluate our HiLo on in-the-wild images, we follow ICON to collect 200 diverse images from Pinterest [^2] https://www.pinterest.com/. The images contain humans performing dramatic movements or wearing diverse clothes.
Figure 6: Visualization results of 3D clothed avatar reconstruction with our HiLo from in-the-wild images, which present various clothing and challenging poses. We show the front (blue) and rotated (red) views.
Table 1: Toy experiments about the impact of SDF on 3D clothed human reconstruction, on seen (i.e., training and test on the same dataset) and unseen (i.e., training on Thuman2.0 and test on CAPE) settings.
Dataset CAPE-FP CAPE-NFP CAPE Thuman2 CAPE
Methods Chamfer ($\downarrow$) P2S ($\downarrow$) Normals ($\downarrow$) Chamfer P2S Normals Chamfer P2S Normals Chamfer P2S Normals Chamfer P2S Normals
$\mathrm{PIFu}$ [41] 2.1000 2.0930 0.0910 2.9730 2.9400 0.1110 2.6820 2.6580 0.1040 2.6880 2.5730 0.0970 7.1244 2.7633 0.3902
$\mathrm{PIFu}_{\mathrm{w}~{}\mathrm{SDF}}$ 0.8908 0.8637 0.0676 0.9848 0.9545 0.0698 0.9437 0.9178 0.0707 1.6659 1.7934 0.1360 1.3078 1.4306 0.0980
$\mathrm{PaMIR}$ [60] 1.2250 1.2060 0.0550 1.4130 1.3210 0.0630 1.3500 1.2830 0.0600 1.4388 1.5613 0.1071 0.9339 0.9444 0.0659
$\mathrm{PaMIR}_{\mathrm{w}~{}\mathrm{SDF}}$ 0.9188 0.8788 0.0565 1.1132 1.0729 0.0611 1.0112 0.9725 0.0601 1.4073 1.5624 0.1174 0.8438 0.8179 0.0572
ICON [50] 0.7475 0.7488 0.0508 0.8656 0.8639 0.0545 0.8055 0.8084 0.0539 1.1431 1.3020 0.0923 0.8610 0.8878 0.0606
$\mathrm{ICON}_{\mathrm{w/o}~{}\mathrm{SDF}}$ 1.0243 0.9478 0.0741 1.4862 1.3313 0.0919 1.2736 1.1538 0.0850 1.3114 1.2116 0.1015 7.4892 1.7708 0.3780
Table 2: (A) Comparison experiments and (B) ablation studies on seen (i.e., training and test on the same dataset) and unseen (i.e., training on Thuman2.0 and test on CAPE) settings. The bold and the underlined numbers indicate the best and second-best results, respectively. "-" denotes that PIFuHD and ECON does not provide a training code.
Train on Thuman2.0 and test on CAPE Train and test on the same dataset (Thuman2.0 or CAPE)
Group Dataset CAPE-FP CAPE-NFP CAPE Thuman2.0 CAPE
Methods Chamfer ($\downarrow$) P2S ($\downarrow$) Normals ($\downarrow$) Chamfer P2S Normals Chamfer P2S Normals Chamfer P2S Normals Chamfer P2S Normals
A PIFu [41] 2.1000 2.0930 0.0910 2.9730 2.9400 0.1110 2.6820 2.6580 0.1040 2.6880 2.5730 0.0970 7.1244 2.7633 0.3902
PIFuHD [42] 2.3020 2.3350 0.0900 3.7040 3.5170 0.1230 3.2370 3.1230 0.1120 2.4613 2.3605 0.0924 - - -
PaMIR [60] 1.2250 1.2060 0.0550 1.4130 1.3210 0.0630 1.3500 1.2830 0.0600 1.4388 1.5613 0.1071 0.9339 0.9444 0.0659
ICON [50] 0.7475 0.7488 0.0508 0.8656 0.8639 0.0545 0.8055 0.8084 0.0539 1.1431 1.3020 0.0923 0.8610 0.8878 0.0606
ECON [51] 0.9651 0.9175 0.0412 0.9983 0.9694 0.0415 0.9872 0.9521 0.0414 - - - - - -
D-IF [52] 0.8038 0.7766 0.0546 0.9877 0.9491 0.0611 0.8878 0.8574 0.0589 1.0305 1.0864 0.0830 0.8332 0.8489 0.0597
B $\mathrm{HiLo}_{\mathrm{w/o}~{}\mathcal{H}(s;\beta)}$ 0.7564 0.7449 0.0514 0.8697 0.8658 0.0553 0.8118 0.8045 0.0547 0.9442 1.0323 0.0785 0.7971 0.7999 0.0551
$\mathrm{HiLo}_{\mathrm{w/o}~{}\mathcal{\phi}_{si}}$ 0.7996 0.7860 0.0569 0.9112 0.9042 0.0601 0.8555 0.8449 0.0468 0.9886 1.0836 0.0850 0.7999 0.7948 0.0547
$\mathrm{HiLo}_{\mathrm{w/o}~{}\mathcal{H}(s;\beta)~{}\mathrm{w/o}~{}\mathcal{\phi}_{si}}$ 0.8662 0.8970 0.0647 1.0201 1.0525 0.0706 0.9362 0.9720 0.0690 1.1220 1.2544 0.0954 0.8125 0.8224 0.0588
HiLo (Ours) 0.6954 0.6876 0.0471 0.7830 0.7876 0.0499 0.7430 0.7428 0.0499 0.9230 0.9855 0.0732 0.7861 0.7729 0.0544
Table 3: Toy experiments on 3D reconstruction with different levels of SMPL-X noise in terms of chamfer distance on unseen CAPE dataset. The voxel grid of naked body $\mathcal{M}_{v}^{3D}(p)$ improves the robustness of reconstruction.
SMPL-X Noise=0.1 SMPL-X Noise=0.2 SMPL-X Noise=0.5
Methods $\mathcal{M}_{v}^{3D}(p)$ CAPE-FP CAPE-NFP CAPE CAPE-FP CAPE-NFP CAPE CAPE-FP CAPE-NFP CAPE
ICON [50] 1.7949 2.0537 1.9284 2.6695 3.0917 2.9365 4.2181 4.8266 4.6716
$\mathrm{ICON}_{\mathrm{w}~{}\mathcal{M}_{v}^{3D}(p)}$ 1.4381 1.5380 1.3411 2.1950 2.3067 2.0723 2.2129 2.3760 2.1188
$\mathrm{D\text{-}IF}$ [52] 1.6078 1.7881 1.7037 2.6853 3.1002 2.9962 4.3591 4.8006 4.7310
$\mathrm{D\text{-}IF}_{\mathrm{w}~{}\mathcal{M}_{v}^{3D}(p)}$ 1.3557 1.7877 1.6064 1.8022 2.4110 2.1959 3.0307 4.4190 4.0807
$\mathrm{HiLo}_{\mathrm{w/o}~{}\mathcal{M}_{v}^{3D}(p)}$ 1.9518 2.1966 2.0877 2.9315 3.4561 3.2661 4.6382 5.0606 4.9844
$\mathrm{HiLo}$ 1.0517 1.3210 1.2004 1.0893 1.5876 1.3427 1.0960 1.6156 1.3593
Table 4: Ablations of our progressive high-frequency SDF on 3D reconstruction with different levels of SMPL-X noise in terms of chamfer distance on unseen CAPE dataset. In addition to $\mathcal{M}_{v}^{3D}$, our progressive high-frequency SDF is also to handle SMPL-X noise due to the coarse-to-fine learning manner.
SMPL-X Noise=0.1 SMPL-X Noise=0.2 SMPL-X Noise=0.5
Methods $\mathcal{M}_{v}^{3D}(p)$ $\mathcal{H}_{s}(p;\beta)$ $\mathcal{H}_{s}(p)$ $SDF$ CAPE-FP CAPE-NFP CAPE CAPE-FP CAPE-NFP CAPE CAPE-FP CAPE-NFP CAPE
$\mathrm{HiLo}_{\mathrm{w/o}~{}\mathcal{H}_{s}(p;\beta)}$ 1.1435 1.4700 1.3124 1.3401 1.9339 1.6909 1.2861 1.8200 1.5620
$\mathrm{HiLo}_{\mathrm{w}~{}\mathcal{H}_{s}(p)}$ 1.1932 1.5541 1.3904 1.1243 1.5575 1.3701 1.2794 1.8973 1.6652
$\mathrm{HiLo}$ 1.0517 1.3210 1.2004 1.0893 1.5876 1.3427 1.0960 1.6156 1.3593
Metrics: We evaluate our HiLo and baseline methods in terms of three metrics: Chamfer distance and P2S distance mainly measure coarse geometry error, while Normals mainly captures high-frequency differences. See details in Appx. C.2.
Baselines: We compare our proposed HiLo with mainstream state-of-the-art methods, including PIFu [41], PaMIR [60], ICON [50], ECON [51] and D-IF [52], refer to the Appx. C.4 for the detailed description. To demonstrate the necessity of naked 3D body regularization, we first conduct a toy experiment to study the effect of SDF on different baselines. To this end, we design three variants based on PIFu, PaMIR, and ICON, which incorporate SDF into existing methods (PIFu and PaMIR) or remove SDF from the existing method (ICON), namely $\mathrm{PIFu}_{\mathrm{w}~{}\mathrm{SDF}}$, $\mathrm{PaMIR}_{\mathrm{w}~{}\mathrm{SDF}}$ and $\mathrm{ICON}_{{w/o}~{}\mathrm{SDF}}$, respectively.
For ablation studies, we construct several variants of our HiLo: 1) $\mathrm{HiLo}_{{w/o}~{}\phi_{si}}$ replaces our spatial interaction implicit function with the vanilla implicit function; 2) $\mathrm{HiLo}_{{w/o}~{}\mathcal{M}_{v}^{3D}}$ is constructed by removing the voxelized SMPL-X; 3) $\mathrm{HiLo}_{{w/o}~{}\mathcal{H}(s;\beta)}$ is constructed by replacing our progressive HF SDF with vanilla SDF.
Implementation details: We implement our approach using PyTorch [^3] We will release our code. [37] and train our networks with RMSprop [47] optimizer. For a fair comparison, we follow all common hyper-parameter settings same as ICON [50]. See more implementation details in the Appx.
Figure 7: Reconstruction results with or without our progressive high-frequency SDF $\mathcal{H}(s;\beta)$. The geometry details of clothes, hands, faces are enhanced by introducing $\mathcal{H}(s;\beta)$.
4.1 Toy Experiments
Our motivation stems from the idea that HF information and LF information improve details and robustness, respectively. To verify this, we employ two tools, i.e.,  SDF and voxelized SMPL-X $\mathcal{M}_{v}^{3D}$ to establish this constraint.
Impact of SDF. We build upon three representative methods, i.e., PIFu, PaMIR, and ICON for the experiments. Specifically, $\mathrm{PIFu}_{{w}~{}\mathrm{SDF}}$ adds SDF following equations: $\mathcal{\phi}(\mathcal{F}_{s}(\mathbf{p}),f_{\text{2D}}(\mathcal{I})({\mathbf{p}}))\rightarrow\hat{\mathcal{O}}$. $\mathrm{PaMIR}_{{w}~{}\mathrm{SDF}}$ incoorporates SDF following: $\mathcal{\phi}(\mathcal{F}_{s}(\mathbf{p}),f_{\text{2D}}(\mathcal{I})({\mathbf{p}}),\mathcal{V}(\mathcal{M})({\mathbf{p}}))\rightarrow\hat{\mathcal{O}}$. $\mathrm{ICON}_{{w/o}~{}\mathrm{SDF}}$ replaces SDF with the z coordinate of $\mathbf{p}$ following: $\mathcal{\phi}(\mathcal{F}_{\mathrm{n}}^{\mathrm{b}}(\mathbf{p}),\mathcal{F}_{\mathrm{n}}^{\mathrm{c}}(\mathbf{p}),\mathbf{p}_{z})\rightarrow\hat{\mathcal{O}}$, where $\mathcal{\phi}$ is the vanilla implicit function and $f_{2D}$ is a 2D CNN. Experimental results in Fig. 2, and Tab. 1 demonstrate the SDF improves geometry details compared with variant methods without it.
Impact of voxelized SMPL-X $M_{v}^{3D}$. Our empirical verification is based on ICON, D-IF and HiLo that requires SMPL-X for reconstruction. we design three variants named $\mathrm{ICON}_{\mathrm{w}~{}\mathcal{M}_{v}^{3D}(p)}$, $\mathrm{D}$-$\mathrm{IF}_{\mathrm{w}~{}\mathcal{M}_{v}^{3D}(p)}$ and $\mathrm{HiLo}_{\mathrm{w/o}~{}\mathcal{M}_{v}^{3D}(p)}$ that add or remove $M_{v}^{3D}$. From the experimental results in Fig. 2 and Tab. 3, we find that incorporating $\mathcal{M}_{v}^{3D}$ helps achieve a more robust reconstruction even faces various levels of noise in SMPL-X shape and pose. The reason is that the low-resolution voxel grid of the naked body is insensitive to noise, and therefore provides robust low-frequency regularization in the training process.
Figure 8: Reconstructions w and w/o our spatial interaction implicit function $\phi_{si}$. Our $\phi_{si}$ is able to perceive the global human body and is therefore able to remove non-human shapes.
4.2 Comparison Experiments
Quantitative results. We conducted comparative experiments in Tab. 2 under two settings. 1) Setting1: Following the setting of the previous methods, we train and test on the same datasets. 2) Setting2: To further evaluate the generalization ability of our HiLo on unseen datasets, we train and test HiLo using different datasets. Our approach achieves the best results in the seen and the unseen settings due to the high- and low-frequency paradigm.
Figure 9: Visualization comparisons on CAPE dataset. The model is training on Thuman2.0 dataset.
Visualization Results. We compare our HiLo with baselines on in-the-wild images and CAPE dataset in Fig. 1, and Fig. 9, respectively. The results show that our HiLo is able to reconstruct 3D clothed avatars with more realistic details. Although ECON obtains detailed fingers by replacing the hand of the SMPL-X model, there exists misalignment on the connection wrist when the corresponding SMPL-X is inaccurate. We put more visualization results of our HiLo on in-the-wild images in Fig. 6. The results demonstrate the effectiveness and generalization ability of our HiLo in recovering detailed geometry (such as hairs and cloth wrinkles). We put more visualization results in the Appx.
4.3 Ablation Studies
How does $\mathbf{\mathcal{H}(s;\beta)}$ improve geometry details? We quantitatively demonstrate the necessity of $\mathcal{H}(s;\beta)$ in Tab. 2. The results demonstrate that $\mathrm{HiLo}_{{w/o}~{}\mathcal{H}(s;\beta)}$ , the variant method that replaces $\mathcal{H}(s;\beta)$ with standard SDF, achieves inferior performance than HiLo. To further study the impact of $\mathcal{H}(s;\beta)$ on common (-FP) and challenging poses (-NFP), we evaluate HiLo on cape dataset that contains both categories. Tab. 2 demonstrates that $\mathcal{H}(s;\beta)$ improves the performance of avatar reconstruction more on challenging poses ($9.54\%$ improvement) than in fashion ($5.21\%$ improvement in terms of Chamfer distance. Furthermore, Fig. 7 demonstrates that $\mathcal{H}(s;\beta)$ leads to more detailed reconstruction, resulting in clearer arms and more realistic cloth wrinkles. From the results, we observe that incorporating the power of high frequency with SDF helps in capturing detailed geometry.
How does $\mathbf{\phi_{si}}$ improve body topology of the reconstructed avatar? As shown in Fig. 8 and Tab. 2, our $\phi_{si}$ removes the non-human shape and boosts reconstruction performance. The reason is that our $\phi_{si}$ leverages a cross-scale attention module $\mathcal{A}$ that builds topological signals between different spatial points in the body model.
4.4 Further Discussions
Is our $\mathbf{\mathcal{H}_{s}(p;\beta)}$ able to help HiLo be robust to SMPL-X noise? We study the impact of $\mathcal{H}_{s}(p;\beta)$ on the robustness ability of our HiLo by replacing it with conventional high frequency SDF $\mathcal{H}_{s}(p)$ and vanilla SDF. We perturb the SMPL-X model with various levels of noise to compare the robustness of the SDF variants. As illustrated in Tab. 4, our proposed $\mathcal{H}_{s}(p;\beta)$ outperforms SDF and high-frequency SDF variants under multiple noise scales due to the progressive manner. See more results in the Appx.
Is HiLo able to converge faster?
Figure 10: Convergence curves of different methods on CAPE dataset. Our HiLo is able to converge faster than existing methods.
In the comparison of validation accuracy depicted in Fig. 10, it is evident that our HiLo  exhibits a remarkable ability to rapidly converge and attain superior performance. Specifically, HiLo swiftly reaches a commendable accuracy of 0.90 at approximately iteration 100. In contrast, the second best method, i.e., ICON, takes significantly longer, around iteration 200, to reach the same level of accuracy, underscoring the efficiency and efficacy of our approach.
5 Conclusion
In this paper, we propose a high-frequency and low-frequency paradigm by exploiting high-frequency and low-frequency information from parametric body models. Based on the paradigm, we design clothed human reconstruction with high- and low-frequency information, namely HiLo that contains: 1) a progressive high-frequency SDF to improve geometry details and alleviate large gradients that hinder model convergence; 2) a spatial interaction implicit function that utilizes the low-frequency complementary information from the voxelized naked body to improve robustness against noise. Experimental results demonstrate the superiority of our HiLo. In the future, we will apply our method to more 3D reconstruction tasks such as 3D face reconstruction, and indoor scene 3D reconstruction.
6 Acknowledgement
This work was partially supported by National Natural Science Foundation of China (NSFC) 62072190, Program for Guangdong Introducing Innovative and Entrepreneurial Teams 2017ZT07X183, and TCL Science and Technology Innovation Fund.
Supplementary Material
[…]
A Related Works
Explicit-shape-based approaches rely on parametric human body models, e.g., SCAPE [4], SMPL [26], SMPL-X [38] to reconstruct 3D humans. Many works [1, 2, 28, 49, 61, 7] introduce the concept of "body+offset", where clothing geometry is represented as 3D displacements on top of the SMPL models. For example, MGN [7] proposes a top-down objective function to align the segmentation maps of predicted garments and SMPL. To improve the expression ability of garment templates and support more topologies, BCNet [22] disentangles the skinning weight of the garment from the body mesh. Different from the representation of "body+offset", alternative parametric methods adapt vertex deformations on body mesh to capture cloth details. For example, HMD [61] presents the hierarchical deformation framework to recover a detailed human body shape from an initial SMPL mesh in a coarse-to-fine manner. The advantage of these methods lies in their compatibility with the current animation pipeline and ease of control through pose parameters. However, they have limitations in modeling various and complex clothing topologies due to the inherent topology constraints imposed by parametric models.
Implicit-function-based approaches aims to reconstruct detailed surfaces with arbitrary topology [12, 29, 34]. This is achieved through the implicit functions, which can be used to approximate 3D representation such as occupancy fields or signed distance fields. PIFu [41] is the pioneering method that utilizes pixel-aligned features for the regression of the occupancy field of human shape. PIFuHD [42] incorporates a multi-level architecture and additional normals to improve the geometric details of PIFu. However, these two methods lack constraints on the global topology of humans, leading to performance degradation in challenging poses. Many works attempt to address this issue in different ways, such as introducing a coarse shape of volumetric humans [18], leveraging depth information of RGB-D images [14]. Unlike the above methods, alternative implicit-function-based methods learn the latent representation of clothing to control the generation of clothing [31, 25, 13]. For example, SMPLicit [13] reconstructs the clothed human by optimizing the latent space of the clothing model to control clothing cut and style. However, the reconstructed human still does not align well with the input image and lacks geometric details.
Explicit shape & Implicit function approaches leverage human body models and implicit functions to harness the benefits of both worlds [8, 9, 20]. For instance, PaMIR [60] regularizes the free-form implicit function by incorporating semantic features from the SMPL model. ICON[50], on the other hand, regresses shapes from locally queried features to generalize to unseen poses in in-the-wild photos. ECON[51] combines estimated 2.5D front and back surfaces with an underlying 3D parametric body for improved reconstruction. To further address the variations in distribution among different spatial points, D-IF[52] introduces a distribution to express the uncertainty of clothing. However, these approaches may fall short when performing highly detailed and robust reconstruction. Specifically, PaMIR is sensitive to global pose and lacks robustness to unseen poses [50]. ECON is prone to reconstruct combined or broken limbs due to the need to complete different surfaces. ICON and D-IF tend to fail in reconstructing detailed parts such as elbows and wrinkles in clothes (see Fig. 1.To promote a detailed reconstruction, our HiLo  uses a progressive high-frequency function to improve the expression of reconstructed details. At the same time, HiLo  uses the low-frequency-based spatial interactive implicit function to enhance robustness to unseen shapes and poses.
B More Details of our HiLo
B.1 Novelty and differences from previous methods  [18, 42, 50, 60]
Clothed human reconstruction from a single RGB image is challenging due to limited views and the absence of depth information. Most recent methods  [50, 51, 60, 52] rely on parametric body models estimated from RGB images, but they may incur an oversmooth problem due to the underutilization of high-frequency (HF) details. Moreover, these methods can be sensitive to noises incurred by parametric body model estimation for challenging poses.
To address the above issues, we first enhance HF information from the body models to describe geometry details. To this end, we design a progressive growing function to achieve accurate reconstruction while alleviating the convergence difficulty associated with HF information. Moreover, we verify low-frequency (LF) information from the parametric model is insensitive to noise. Considering this, we establish a spatial interaction function to leverage the (LF) for robustness reconstruction.
B.2 Details on Spatial Interaction MLP
Figure A: Illustration of the spatial interaction module $\mathcal{A}$.
Our spatial interaction implicit function takes $\mathbf{F}_{c}^{1}$ that contains our high-frequency SDF $\mathcal{H}(s;\beta)$, low-frequency voxel grids feature $\mathcal{M}^{3D}_{\mathcal{V}}({\mathbf{p}})$, and normal features $\mathbf{F}_{\mathrm{n}}(\mathbf{p})$ as input and infers occupancy fields $\mathcal{\hat{O}}$.
$\mathbf{F}_{c}^{1}=[\mathcal{H}(s;\beta),\mathcal{M}^{3D}_{\mathcal{V}}({\mathbf{p}}),\mathbf{F}_{\mathrm{n}}(\mathbf{p})]$ (A)
${\phi_{si}}(\mathbf{F}_{c}^{1})\to\mathcal{\hat{O}},~{}~{}\phi_{si}(\cdot)=\mathcal{A}^{N+1}\circ T^{(N+1)}\circ\cdots\circ\mathcal{A}^{1}(\cdot)\circ T^{(1)}.$ (B)
As shown in Fig. A, take the $1$-th layer of $\phi_{si}$ as an example, the attention module $\mathcal{A}^{1}$ takes in the $\mathcal{M}^{3D}_{\mathcal{V}}$ and $\mathbf{F}_{c}^{(1)}$ and output a spatial interaction feature map $\mathbf{F}_{I}^{(1)}$. Specifically, We first extract a global spatial features $\mathbf{F}_{g}^{(j)}$ of the $\mathcal{M}_{v}^{3D}$ via a 3D Convolution block and a sigmoid function. We achieve the spatial interaction process of different voxels through the equation $\mathbf{F}_{c}^{(j)}\times\mathbf{F}_{g}^{(j)}\rightarrow\mathbf{F}_{I}^{(j)}$. After obtaining the $\mathbf{F}_{I}^{(1)}$, we fed it to the first full-connected layer $T^{(1)}$ to obtain $\mathbf{F}^{2}_{c}$.
B.3 Future work and limitations.
Q5. Future work and limitations. Since HiLo is trained on orthographic views, it struggles with strong perspectives, causing asymmetrical limbs or unrealistic shapes. This issue is worth studying in the future.
C More Experimental Details
Input: Sampled 3D points $\{\mathbf{p}\}_{i=1}^{n}$, an RGB image $\mathcal{I}$ of human, an spatial interactive implicit function $\phi_{si}$, a parametric body model estimation net $E_{p}$, a progressive high frequency function $\mathcal{H}(\cdot~{};\beta)$, a 3D CNN $f_{3D}$, a mesh voxelization operation $\mathcal{V}$, a marching cubes operation $\mathcal{MC}$.
Output: Triangular mesh of the human.
1 Obtaining parametric body model SMPL-X $\mathcal{M}$ with $E_{p}(\mathcal{I})$.
2 With $\mathcal{M}$, obtaining the global voxel grid $\mathcal{M}_{v}^{3D}$ using $f_{3D}(\mathcal{V}(\mathcal{M}))$.
3 for $\mathbf{p}_{i}$ in $\{\mathbf{p}\}_{i=1}^{n}$ do
4       Generating SDF $s$ w.r.t. $\mathbf{p}_{i}$ using Eqn. 1.Using $\mathcal{H}(\cdot~{};\beta)$ to enhance the SDF $s$ resulting in point-wise progressive high-frequency SDF $\mathcal{H}(s;\beta)$.
5       Obtaining the local voxel grid of ${\mathcal{V}}$ by indexing $\mathcal{M}_{v}^{3D}$ with $\mathbf{p}_{i}$, resulting in ${\mathcal{V}}(\mathbf{p}_{i})$.
6       Get 3D normal features $\mathbf{F}_{\mathrm{n}}(\mathbf{p}_{i})$ w.r.t. $\mathbf{p}_{i}$ following ICON.
7       Concatenate $\mathcal{H}(s;\beta)$, ${\mathcal{V}}(\mathbf{p})$, $\mathbf{F}_{\mathrm{n}}(\mathbf{p}_{i})$, getting $F_{c}^{1}$.
8       Using $\phi_{si}$ to obtaining occupancy field $\hat{{\mathcal{O}}}(\mathbf{p}_{i})$ from $F_{c}^{1}$ and $\mathcal{M}_{v}^{3D}$, following Eqn. (7).
9 end for
Obtaining the triangular mesh of the human using marching cubes algorithm with $\mathcal{MC}(\hat{{\mathcal{O}}})$.
Algorithm 1 The inference pipeline of HiLo.
We demonstrate the inference details of our HiLo in Alg. 1. The 3D point set is obtained via a coarse-to-fine manner as illustrated in Sec. C.3.
C.1 Implementation Details
Especially, the dimension of HFSDF, batch size $\mathrm{b}$, sampled points number $\mathrm{n}$, and variable dimension channels $\mathrm{C}$ of the spatial interaction module are set to $10$, $2$, $8000$, $[39,512,256,128,1]$ respectively. The training and testing phases are performed on a single NVIDIA GeForce RTX 3090 GPU. See more details on the training and inference of HiLo in the appendix.
C.2 More details on Metrics
Specifically, P2S denotes the distance between randomly sampled points from a ground truth mesh to its nearest surface on a reconstructed mesh. Chamfer is regarded as a bidirectional P2S distance, which computes the distance between randomly sampled points from the reconstructed mesh to its nearest surface on the ground truth mesh. Normals is calculated by measuring L2 error between normal images rendered from reconstructed and ground-truth meshes from fixed viewpoints.
C.3 3D Points Sampling
During training, we randomly query 3D points inside, outside, and around the SMPL-X surface. During inference, we define the coordinates of 3D points through an initial 3D grid, and iteratively interpolate the 3D grid to sample 3D points in a more detailed scale.
C.4 Details of Variant Methods
C.4.1 Revisit Existing Methods
PIFu. To reconstruct a 3D-clothed human, PIFu proposes Pixel-Aligned Implicit Functions to predict whether each 3D point is inside or outside a human surface. Specifically, PIFu learns a 2D feature map from a single image $I$ using a 2D image encoder via $f_{\text{2D}}(\mathcal{I})\rightarrow\mathcal{F}_{I}^{2D}$. To query local pixel-aligned features on $\mathcal{F}_{I}^{2D}$, PIFu projects 3D points $\mathbf{p}$ to a 2D plane with $\pi$ operation and uses bilinear interpolation operation $\mathrm{S}$ to sample the local features from $\mathcal{F}_{I}^{2D}$. The local feature $f_{\text{2D}}(\mathcal{I})(\mathbf{p})$ and the Z coordinate of $\mathbf{p}$ (i.e.,  $\mathbf{p}_{z}$) are concatenated and fed to a multi-layer perceptron (MLP) to obtain the final prediction $\hat{\mathcal{O}}$. The pipeline of PIFU follows an equation:
$\mathrm{PIFu}\mathrel{\mathop{:}}\phi(f_{\text{2D}}(\mathcal{I})(\mathbf{p})),\mathbf{p}_{z})\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$ (C)
where $f_{\text{2D}}$ denotes the 2D image encoder. Although PIFu is able to reconstruct high-quality human mesh for commonly seen poses such as walking and standing, PIFu often fails when encountering severe occlusions and large pose variations due to insufficient information from a single image only.
PaMIR. To further regulate the reconstruction process, PaMIR introduces the strengths of parametric body models by learning a parametric-aligned 3D feature volume acquired from a parametric body model, i.e.,  SMPL. Specifically, PaMIR estimates a SMPL model $\mathcal{M}$ from the given single image $I$, converting $\mathcal{M}$ to occupancy volume with mesh voxelization $\mathcal{V}$ and encoding the volume with 3D convolutional neural networks $f_{\text{3D}}$. Given the voxel-aligned volume features $f_{\text{3D}}(\mathcal{V}(\mathcal{M})),{\mathbf{p}}))$ and the corresponding pixel-aligned feature vector $f_{\text{2D}}(\mathcal{I})({\mathbf{p}}))$ of $\mathbf{p}$, PaMIR learns an implicit function to predict whether $\mathbf{p}$ is inside or outside a human surface. The pipeline of PaMIR follows the equation:
$\mathrm{PaMIR}\mathrel{\mathop{:}}\phi((f_{\text{2D}}(\mathcal{I})({\mathbf{p}})),\mathcal{M}_{v}^{3D}({\mathbf{p}}))\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$ (D)
Although PaMIR typically feeds their implicit-function module with features of a global 2D image or 3D voxel encoder, but these features are sensitive to global pose [50].
ICON. To improve the robustness to out-of-distribution poses, ICON replaces the global encoder of existing methods with a more local scheme: using signed distance function (SDF), barycentric surface normal and local normal features of SMPL regarding $\mathbf{p}$. The pipeline of ICON follows the equation:
$\mathrm{ICON}\mathrel{\mathop{:}}\phi(s(\mathbf{p}),\mathcal{F}_{\mathrm{n}})\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$ (E)
where $\mathcal{F}_{s}(\mathbf{p})$ is the signed distance from a query point $\mathbf{p}$ to the closest body point $\text{P}^{\text{b}}\in\mathcal{M}$, and $\mathcal{F}_{\text{n}}^{\text{b}}$ is the barycentric surface normal of $\text{P}^{\text{b}}$, and $\mathcal{F}_{\text{n}}^{\text{c}}$ is a normal vector. We denote the concatenation of $\mathcal{F}_{\mathrm{n}}^{\mathrm{b}}(\mathbf{p})$, $\mathcal{F}_{\mathrm{n}}^{\mathrm{c}}(\mathbf{p})$ as $\mathcal{F}_{\mathrm{n}}$.
D-IF. To alleviate the uncertainty in the process of reconstructing a clothed human, D-IF follows ICON to estimate the occupancy field of the clothed human based on the equation:
$\mathrm{D\text{-}IF}\mathrel{\mathop{:}}~{}$ $\hat{{\mathcal{O}}}_{f}=\hat{{\mathcal{O}}}_{c}+\phi_{r}(\hat{{\mathcal{O}}}_{c}\oplus{\mathcal{F}}_{7\mathrm{D}}\oplus P_{\varphi}({\mathcal{F}}_{7\mathrm{D}}))$ (F)
${\mathcal{F}}_{7D}=s\oplus\mathcal{F}_{\mathrm{n}},~{}\hat{{\mathcal{O}}}_{c}=\phi({\mathcal{F}}_{7\mathrm{D}})$
$\hat{{\mathcal{O}}}_{c}(p)\sim P_{\varphi}\left(F_{7\mathrm{D}}({\mathbf{p}})\right)=\mathcal{N}\left(\mu_{\varphi}({\mathbf{p}}),\sigma_{\varphi}({\mathbf{p}})\right)$
where $\oplus$ denotes concatenate operation, $P_{\varphi}(F_{7\mathrm{D}}({\mathbf{p}})$ is a Gaussian distribution.
C.5 Variant Methods
Based on the grasp of existing methods, we introduce the variant methods in our experiments.
$\mathrm{ICON}_{\mathrm{w}~{}\mathcal{M}_{v}^{3D}(p)}\mathrel{\mathop{:}}\phi(s({\mathbf{p}}),\mathcal{F}_{\mathrm{n}},\mathcal{M}_{v}^{3D}({\mathbf{p}}))\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$ (G)
$\mathrm{D\text{-}IF}_{\mathrm{w}~{}\mathcal{M}_{v}^{3D}(p)}\mathrel{\mathop{:}}\phi({\mathcal{F}}_{7\mathrm{D}},\mathcal{M}_{v}^{3D}({\mathbf{p}}))+\phi_{r}(\hat{{\mathcal{O}}}_{c}\oplus{\mathcal{F}}_{7\mathrm{D}}\oplus P_{\varphi}({\mathcal{F}}_{7\mathrm{D}}))$
$\mathrm{HiLo}_{\mathrm{w/o}~{}\mathcal{M}_{v}^{3D}(p)}\mathrel{\mathop{:}}\phi_{si}(\mathcal{H}(s;\beta),\mathcal{F}_{\mathrm{n}})\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$
$\mathrm{HiLo}_{\mathrm{w/o}~{}\mathcal{H}_{s}(p;\beta)}\mathrel{\mathop{:}}\phi_{si}(s({\mathbf{p}}),\mathcal{F}_{\mathrm{n}},\mathcal{M}_{v}^{3D}({\mathbf{p}}))\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$
$\mathrm{HiLo}_{\mathrm{w}~{}\mathcal{H}_{s}(p)}\mathrel{\mathop{:}}\phi_{si}({\mathcal{H}}(s),\mathcal{F}_{\mathrm{n}},\mathcal{M}_{v}^{3D}({\mathbf{p}}))\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$
$\mathrm{HiLo}_{\mathrm{w/o}~{}\mathcal{\phi}_{si}}\mathrel{\mathop{:}}\phi(\mathcal{H}(s;\beta),\mathcal{F}_{\mathrm{n}},\mathcal{M}_{v}^{3D}({\mathbf{p}}))\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$
$\mathrm{HiLo}_{\mathrm{w/o}~{}\mathcal{H}(s;\beta)~{}\mathrm{w/o}~{}\mathcal{\phi}_{si}}\mathrel{\mathop{:}}\phi(s({\mathbf{p}}),\mathcal{F}_{\mathrm{n}},\mathcal{M}_{v}^{3D}({\mathbf{p}}))\rightarrow\hat{\mathcal{O}}(\mathbf{p}))$
D More Details on Datasets
Our data-split configuration aligns with the protocols outlined by ICON and D-IF. We conduct experiments on the basis of two distinct settings.
  • Setting 1: Train on Thuman2.0, test on CAPE. For this setting, we employ 500 scans from Thuman2.0 for training, accompanied by 5 scans for validation. To assess reconstruction accuracy on CAPE, we utilize 150 scans, further categorized into challenging poses ("CAPE-NFP" - 100 scans) and fashion poses ("CAPE-FP" - 50 scans). To emulate diverse viewpoints during testing, RGB images are synthesized by rotating a virtual camera around the textured scans at angles of $0^{\circ},120^{\circ},240^{\circ}$.
  • Setting 2: Train and test on the same dataset. In this scenario, when training and testing on Thuman2.0, we employ 500 scans for training and reserve 20 scans for testing. Conversely, when training and testing on CAPE, we utilize 120 scans for training, 5 for validation, and 25 for testing.
E More Experiments
E.1 More results on SMPL-X noise.
SMPL-X Model. Skinned Multi-Person Linear-Expressive model (SMPL-X) [38] represents human body shapes and poses in a compact and parametric manner. The core idea behind SMPL is to use a linear combination of body shape parameters and joint rotations to represent a 3D human body model with $N{=}10475$ vertices and $K{=}54$ joints. Specifically, SMPL-X is defined by $\mathcal{M}(\theta,\beta,\psi)\mathrel{\mathop{:}}\mathbb{R}^{|\theta|\times|\beta|\times|\psi|}\to\mathbb{R}^{3N}$, where $\theta\in\mathbb{R}^{3(K+1)}$ represents the pose parameter, $\beta\in\mathbb{R}^{|\theta|}$ is the shape parameter, and $\psi$ denotes the facial expression parameters, and $K$ denotes the number of body joints in addition to a joint for global rotation. By adjusting $\theta$, $\beta$, $\psi$, SMPL-X is able to represent a wide variety of human body shapes and poses. See [38] for more details.
Adding noise to SMPL-X Model. We further evaluate the robustness ability of our HiLo against various levels of noise in the shape parameters $\theta_{s}$ and pose parameters $\theta_{p}$ in parametric models. Our experimental setting follows ICON, which samples a scalar value $\mu~{}\sim\mathcal{N}(0,1)$, scaling the noise with two predefined parameters $s_{1}$, $s_{2}$ to represent various levels of noise. The above procedure follows the equation:
$\theta_{s}+=s_{1}\ast\mu$ (H)
$\theta_{p}+=s_{2}\ast\mu$
We set $\{s_{1},s_{2}\}$ to $\{0.1,0.1\}$, $\{0.2,0.2\}$, $\{0.3,0.3\}$, $\{0.4,0.4\}$, $\{0.5,0.5\}$ for a thorough study on the robustness of our HiLo and our baselines. Since we have provided the results w.r.t.  $\{s_{1},s_{2}\}\in[\{0.1,0.1\},\{0.2,0.2\},\{0.5,0.5\}]$ in the main draft, we report the remaining results in Tab. A
SMPL-X Noise=0.3 SMPL-X Noise=0.4
Methods $\mathcal{M}_{v}^{3D}$ CAPE-FP CAPE-NFP CAPE CAPE-FP CAPE-NFP CAPE
ICON 4.5134 4.7091 4.7069 5.5864 5.9810 5.9015
ICON w $\mathcal{M}_{v}^{3D}$ 4.2250 4.1215 4.2697 3.3824 3.4722 3.3897
D-IF 3.2462 3.6933 3.5700 3.2462 3.6933 3.5700
D-IF w $\mathcal{M}_{v}^{3D}$ 1.2912 1.8222 1.5995 1.2912 1.8222 1.5995
HiLo w/o $\mathcal{M}_{v}^{3D}$ 3.7060 4.3281 4.1071 4.4435 4.8639 4.7763
HiLo 1.1014 1.5407 1.3552 1.1633 1.7584 1.5132
Table A: Impact of $\mathcal{M}_{v}^{3D}$ on different methods in terms of Chamfer Distance. We train the models on Thuman2.0 and test them on CAPE.
E.2 More error measurements to assess robustness.
To further evaluate the robustness of our HiLo, we calculate Chamfer, P2S and Normals between SMPL-X and reconstructed body models. From Tab. E.2, our HiLo shows better robustness than existing methods.
Table B: Robustness on CAPE.
Methods Chamfer ($\downarrow$) P2S ($\downarrow$) Normals ($\downarrow$)
PIFu 4.0550 3.3971 0.1915
PIFuHD 6.1345 5.2692 0.2017
PaMIR 0.9800 1.0132 0.0714
ICON 0.8198 0.7799 0.0617
D-IF 0.9111 0.8751 0.0666
ECON 0.9083 0.8701 0.0723
HiLo (Ours) 0.6784 0.6580 0.0480
E.3 Is HiLo efficient and light-weighted?
Comparison of inference and training time. In Tab. C, we compare the inference efficiency by the average inference time to reconstruct 200 single-view images. The inference procedures of PaMIR, ICON, D-IF, and HiLo consist of SMPL-X fitting and cloth refinement. Differently, PIFu’s inference procedure only includes cloth refinement, and ECON includes SMPL-X fitting and Poisson Surface Reconstruction (PSR). In terms of inference efficiency, it is evident that our HiLo  demonstrates a competitive performance with PaMIR, ICON and D-IF. However, ECON depends on time-consuming PSR to complete human shape, and all other methods show superior performance to it when inference. We measure training efficiency by the average time spent on 10 epochs on the Thuman2 dataset. D-IF needs to train two MLPs and therefore takes more time. We achieve competitive training efficiency with PIFu, PaMIR and ICON even though we introduce high-frequency and low-frequency information simultaneously. ECON lacks this statistic because the authors do not release the training codes. Comparison of model size. Form Tab. C, with the exception of ECON, the model sizes of existing methods are basically the same. Although ECON is lightweight, it requires time-consuming PSR to complete meshes of human shape.
Table C: Comparing training/inference efficiency and model size of existing methods.
Method Inference Time (seconds) Training Time (seconds) Million Parameters (seconds)
PIFu [41] 8.13 1636 28.09
PaMIR [60] 21.97 1298 28.18
ICON [50] 18.63 1697 28.11
D-IF [52] 18.51 2336 28.79
ECON [51] 110.93 - 12.07
HiLo (Ours) 19.17 1918 28.21
F More visualization Results
F.1 Transfer Sketch to 3D model
Since our HiLo is robust to in-the-wild images, we are able to put it to more applications. We show in Fig. B that our HiLo is able to transfer a sketch image of a clothed human into a 3D model with the help of ControlNet [56]. Specifically, we collect sketch images from Pinterest and use ControlNet to transfer the images to RGB images. The RGB images are then fed to our HiLo to reconstruct 3D model of the corresponding human.
Figure B: More application of our HiLo. We are able to transfer a sketch of a clothed human into a 3D model.
F.2 Results on In-the-wild Images
We report more comparisons with state-of-the-art methods on in-the-wild images in Fig. C, Fig. D, Fig. E, Fig. F, Fig. G, Fig. H, Fig. I, Fig. J, Fig. K. We render the reconstructed 3D models from four different views, i.e., $0^{\circ},90^{\circ},180^{\circ},270^{\circ}$.
Figure C: Visualization comparisons of reconstruction for our HiLo  vs SOTA.
Figure D: Visualization comparisons of reconstruction for our HiLo  vs SOTA.
Figure E: Visualization comparisons of reconstruction for our HiLo  vs SOTA.
Figure F: Visualization comparisons of reconstruction for our HiLo  vs SOTA.
Figure G: Visualization comparisons of reconstruction for our HiLo  vs SOTA.
Figure H: Visualization comparisons of reconstruction for our HiLo  vs SOTA.
Figure I: Visualization comparisons of reconstruction for our HiLo  vs SOTA.
Figure J: Visualization comparisons of reconstruction for our HiLo  vs SOTA.
Figure K: Visualization comparisons of reconstruction for our HiLo  vs SOTA.